OpenStudy (anonymous):

Translate the following situation into an equation. Machine A can sort products twice as fast as Machine C and Machine B can sort products at a rate of 300 products fewer per min. than Machine C. If together the three machines have an average sorting rate of 2,150 products per minute, how many products per minute can Machine C sort?

OpenStudy (anonymous):

A=2C , B=C-300 , A+B+C =2150 ===> 2C +C-300+C=2150 ===> 4C= 2450 ===> C=612.5

OpenStudy (anonymous):

this does not have to be solved only the equation written out not sure how to write it though

OpenStudy (anonymous):

which answer is correct

OpenStudy (anonymous):

A=2C [ eq. 1] , B=C-300 [ eq. 2] , A+B+C =2150 [eq. 3] === By subtituting eq. 1 & eq. 2 in eq 3.===> 2C +C-300+C=2150 ===> 4C= 2450 ===> C=612.5

OpenStudy (anonymous):

hope that helps you.

OpenStudy (anonymous):

A=2C,B=C-300,A+B+C/3=2150 put values of A and B in eq(iii) 4C-300=6450 C=1687.5